OpenStudy (anonymous):

Hi guys! Do you know how to solve x(e^x)+1 = 0? Can I use log? Thanks!

OpenStudy (saifoo.khan):

@ChrisS

jhonyy9 (jhonyy9):

yes you need to us ln

jhonyy9 (jhonyy9):

- so because there is e^x

OpenStudy (anonymous):

So ln (x) + ln (e^x) =ln (-1). If we ln both sides. But would that be wrong, since -1 is not in the domain for the natural log function?

jhonyy9 (jhonyy9):

do you know the property of logarithm ? than loga(a) =1 - yes ?

jhonyy9 (jhonyy9):

- and you know that ln is logarithm base e - yes ?

OpenStudy (anonymous):

lhs is always +....

OpenStudy (anonymous):

Well, I moved the +1 to the RHS so that ln of a product would be left on the LHS and thus, I can separate them into products. ln (ab)= ln a + ln b, right? Or am I forgetting something?

OpenStudy (beginnersmind):

"lhs is always +...." True, but how do you prove it? greencat, I subtracted 1 and divided by x. Got e^x=-1/x and graphed both.

OpenStudy (anonymous):

Ahhh. After graphing, how did you get x? :)

OpenStudy (beginnersmind):

Did _you_ try to graph them? :) If there's a solution it's where the two graphs intersect.

OpenStudy (beginnersmind):

To clarify, I don't have a solution. But graphing those two functions gives a pretty good idea of what's going on.

OpenStudy (anonymous):

Look, just differentiate xe^x and go from there.....